OpenStudy (anonymous):

The table below shows the proportional relationship between the weight of yogurt, in ounces, and the number of packs of yogurt: Number of Packs of Yogurt Number of Ounces of Yogurt 2 36 3 54 5 90 What is the constant of proportionality? 10 18 36 57

OpenStudy (anonymous):

help i give fan and medal

OpenStudy (anonymous):

36/2 = 54/3 = 90/5 = 18

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Anytime